# DARTS-GT: Differentiable Architecture Search for Graph Transformers with Quantifiable Instance-Specific Interpretability Analysis

### v1

@@ I Introduction
These discovered architectures must also be interpretable for trustworthiness and scientific validity. Unlike well-developed GNN explainability methods [21] that identify prediction-critical subgraphs, GT interpretation, widely regarded as a blackbox [22], remains dominated by attention visualization [23] lacking quantitative metrics and the ability to differentiate visually prominent but functionally irrelevant patterns from subtle yet critical ones. However, there has been a recent surge in significant GT interpretability efforts that inspect attention patterns (e.g., aggregating head/layer attention into an attention graph) [24, 23]. However, these methods do not test faithfulness (whether heads/nodes causally affect predictions) nor enable cross-model comparison via standardized scores. Task-specific self-explanations exist but do not generalize to arbitrary GTs or provide head attributions [25]. We seek a GT-agnostic instance-level interpretability framework grounded in perturbation/ablation that delivers model-level scores for comparing GT interpretability.
@@
Second, we introduce a Fidelity like [29, 30], instance-level, ablation-grounded score: Head-deviation, quantifying prediction changes when specific heads are masked. From per-head effects on each instance, we compute two complementary metrics: Specialization captures head distinguishability as dispersion of head impacts (high = few heads carry causal load; low = heads act interchangeably); Focus captures consensus over graph regions by measuring top-head agreement on top-k nodes (high = heads converge on same salient substructures or nodes). Aggregating per-instance values by dataset median yields robust model-level summaries. Together, these form our interpretability framework that we propose to analyse our model along with existing models for comparison. Head-deviation provides head and graph-region attributions per input, while median Specialization and Focus enable dataset-level GT comparison indicating which model is easier to interpret at matched accuracy.
@@ III-A3 Feed-Forward Layers and Skip Connections
where $\sigma$ is the nonlinearity (we use gelu in our implementation), $\mathbf{W}_{1}^{(\ell)}\in\mathbb{R}^{d\times rd}$, $\mathbf{W}_{2}^{(\ell)}\in\mathbb{R}^{rd\times d}$, and $\mathbf{b}_{1}^{(\ell)},\mathbf{b}_{2}^{(\ell)}$ are biases. We use r=2r{=}2 in our implementation, and $\mathrm{Norm}$ denotes either LayerNorm or BatchNorm depending on configuration.
@@ III-B DARTS-GT
where $\mathcal{S}$ denotes the search space of all $n^{L}$ architectures, $w$ represents the network weights, and $w^{*}(\mathcal{A})$ are the optimal weights for architecture $\mathcal{A}$.
@@ III-B1 Bilevel Optimization via DARTS
Enumerating all possible configurations in $\mathcal{S}$ is generally infeasible so we adopt the Differential Architecture Search (DARTS) [27] framework to solve the optimization problem formulated by equations 11 and 12. DARTS reformulates discrete selection as a continuous, differentiable optimization. Specifically, it introduces a softmax relaxation over candidate operations, casting the architecture search as the following bilevel optimisation problem [38]:
@@
where $w$ represents the network weights and $\alpha=\{\alpha_{o}^{(\ell)}\mid o\in\mathcal{O},\ell\in\{1,...,L\}\}$ are continuous architecture parameters, with $\alpha_{o}^{(\ell)}\in\mathbb{R}$ representing the architecture weight (importance factor) for the operation $o$ at layer $\ell$. Rather than solving the full bilevel optimization, we adopt the efficient first-order approximation of DARTS [27], which avoids Hessian–vector products. In practice,we first randomly shuffle the entire training dataset $\mathcal{D}_{\text{train}}$ and partition it into darts-training split : $\mathcal{D}_{\text{DARTStrain}}$ for updating weights $w$ and darts-validation split : $\mathcal{D}_{\text{DARTSval}}$ for updating architecture parameters $\alpha$, alternating the single gradient steps [27]. More details about this approach can be found in [27]. Although the second-order DARTS update is more accurate, it requires expensive Hessian–vector products and offers limited practical gains in our setting.
@@ III-B2 Search Space Definition
We define our search space as: $\mathcal{O}$ = {GINE [39], GATv2 [40], GatedGCN [32]}. Hence, our $n=3$ leads to the search space of $3^{L}$ architectures.
These operators were selected because they represent very distinct message-passing paradigms that capture complementary aspects of graph structure:
@@ III-B3 Asymmetric Attention Design
Departing from the standard transformer formalism outlined in the Preliminaries section, the novelty of our approach includes an asymmetric design: queries are derived directly from node features, while keys and values are produced by layer-specific GNN operators. This is initially motivated by the GMH design of GMT [26], where they demonstrated that decoupling the computational sources for the attention components enhances graph-aware representations. In GMT, node features are first transformed by a GNN to obtain $H$, with queries taken directly as $Q=H$ and each head computing its own keys and values via an additional GNN applied to $H$. This asymmetry separates structural encoding from feature representation, but at the cost of computational inefficiency, since every key and value inside every head requires an independent GNN. In contrast, our design takes node features directly into $Q$, and shares the GNN operator across heads within a layer, producing an efficient formulation that still preserves the benefits of feature–structure decoupling. More broadly, rather than augmenting Graph Transformers externally with auxiliary MPNNs, a widely practiced approach [6, 16], we demonstrate that graph awareness can emerge from architectural redesign of the attention mechanism itself instead.
@@
During the search phase, DARTS [27] relaxes the discrete operator selection into a continuous mixture. At each layer $\ell$, instead of selecting a single operator, we compute a weighted combination:
@@
During search, the mixed operator output $\overline{o}^{(\ell)}(\mathbf{Z}^{(\ell)},E)$ provides the structural path for keys and values:
@@
We compute $\overline{o}^{(\ell)}(\mathbf{Z}^{(\ell)},E)$ once per layer and share it across all heads $m$ to form $\mathbf{K}_{m}^{(\ell)}$ and $\mathbf{V}_{m}^{(\ell)}$. After discretization, $\overline{o}^{(\ell)}$ is replaced by the selected operator $o^{(\ell)*}$(see next point). Attention may be dense or sparse as that only changes the softmax normalization (all-pairs vs. edge-softmax) and does not alter the source of $\mathbf{K}_{m},\mathbf{V}_{m}$.
@@ III-B5 Discretization and Final Architecture Training
Following standard DARTS protocol, the final discretized architecture $\mathcal{A}^{*}=\{o^{(1)*},...,o^{(L)*}\}$ is then re-trained from scratch using the full training data. In our implementation, we instantiate a fresh model containing only the selected operations per layer. The algorithm is discussed in Algorithm. 1.
@@ III-B6 Edge Residual Stream
$\mathrm{MLP}_{\text{edge}}:\mathbb{R}^{d_{e}}\!\to\!\mathbb{R}^{d}$ is a two-layer MLP with dropout; $\sigma_{g}(x)$ is sigmoid function and $\gamma^{(\ell)}\!\in\!\mathbb{R}$ is a per-layer learnable scalar controlling the gate; $\mathbf{Y}^{(\ell)}$ is the pre-residual attention output. $\mathrm{MLP}_{\text{edge}}$ is shared across layers and $\mathbf{H}^{\text{edge}}$ is computed once per forward pass. Please see Supp-Table I for ablation studies related to the edge residual stream. This edge residual stream operates independently of the selected GNN operations, allowing the model to leverage edge information even when the discretized architecture uses operations with limited edge feature utilization.
@@ III-B7 Computational Cost
Our proposed asymmetric block applies a single GNN per layer to produce the structural representations used in both $K$ and $V$. In contrast, GMT’s Graph Multi-head Attention (GMH) [26] instantiates per-head GNNs for both keys and values, i.e., $\{\mathrm{GNN}^{K}_{i},\mathrm{GNN}^{V}_{i}\}_{i=1}^{M}$. Consequently, the GNN forward cost in GMH scales as $\mathcal{O}(2M\cdot\text{cost}_{\text{GNN}})$, whereas in our design it is $\mathcal{O}(1\cdot\text{cost}_{\text{GNN}})$ per layer. Thus we preserve feature–structure interaction ($Q$ from features; $K,V$ from GNN) while reducing the structural computation across heads. For the architecture search cost, DARTS [27] uses first-order approximation that avoids the Hessian–vector product overhead.
@@ III-C1 Framework Overview
We propose a systematic framework to quantify Graph Transformer interpretability through head-level and node-level analysis. Our approach goes beyond visualization and yields quantitative, instance-specific evidence of which components causally drive predictions. We introduce Head-Deviation, Specialization, and Focus, aligned with established interpretability principles where fidelity, sparsity, and consensus are standard evaluation criteria [30, 29]. Because these metrics are ablation-grounded, they measure causal contribution rather than correlation, allowing robust, fair comparisons of interpretability across architectures.
@@
For head $m$ in layer $\ell$, we implement a weight-level ablation that removes its contribution at inference. Concretely, we silence head $m$ in layer $\ell$, by zeroing its pre-concatenation output $Y_{m}^{(\ell)}$ (equivalently, zeroing the corresponding slice of $\mathbf{W}_{\text{out}}$ in Eq. 6). All other parameters including the outputs of other heads, layer norms, and feed-forward sublayers are left unchanged. Models are evaluated in eval mode (no dropout), with a fixed seed and without any retraining or adaptation.
@@ III-C2 Head Deviation
To quantify the importance of individual attention heads, we measure how their removal affects the task loss. For a graph $G_{i}$ with true label $y_{i}$, let $f(G_{i})$ denote the model’s prediction and $f^{-m,\ell}(G_{i})$ denote the prediction with head $m$ in layer $\ell$ masked.
@@
Here, $\mathrm{loss}(\cdot,\cdot)$ denotes the dataset-appropriate loss (e.g., MAE for regression, binary cross-entropy for binary classification, cross-entropy for multi-class classification). We quantify head importance through causal intervention: the deviation $\delta_{m,\ell,i}$ measures how the prediction changes when head $m$ in layer $\ell$ is ablated, providing direct evidence of its functional necessity rather than mere correlational patterns. This approach asks ”what happens without this head?” , distinguishes truly important components from those that merely appear salient in visualizations.
@@
For each instance, we rank heads by their deviation $\delta_{m,\ell,i}$ in descending order. Beneficial heads ($\delta_{m,\ell,i}>0$, masking increases loss) thus appear first, while harmful heads ($\delta_{m,\ell,i}<0$) appear last. When we need a sign-agnostic view, we report the magnitude $|\delta_{m,\ell,i}|$ as a measure of head impact.
@@ III-C3 Global Interpretability Metrics
where $\mathrm{std}(.)$ denotes the standard deviation across all heads’ deviations with $\mathcal{M}={[1,...,M]}$ , $\mathcal{L}={[1,...,L]}$ where $M$ is the number of attention-heads and $L$ total number of layers, defined earlier.
High specialization indicates heads are easily distinguishable by their causal impact whereas low specialization suggests similar head impacts, making it harder to isolate drivers unless complemented by the focus metric.
@@
Let $\mathcal{P}_{i}^{*}=\{p_{1},\dots,p_{k}\}$ be the top-$k$ head-layer pairs for instance $G_{i}$, ranked according to the head-ranking rule defined above, where $p_{j}=(m_{j},\ell_{j})$. Here, $m_{j}$, $l_{j}$ (j=1,,kj=1,...,k) are index variables with $m_{j}\in\mathcal{M}$ and $l_{j}\in\mathcal{L}$. For a pair $p$, we define $N_{p}$ as the set of top-rr% nodes by incoming attention mass for that head on $G_{i}$:
  • Dense attention: with head softmax matrix $\mathbf{S}^{(\ell)}_{m}$, incoming mass for node $j$ is $\sum_{i}[\mathbf{S}^{(\ell)}_{m}]_{ij}$; $N_{p}$ contains the top-rr% nodes by this value.
  • Sparse attention: with edge-softmax $s_{m,ij}$ on $(i,j)\in E$, incoming mass for node $j$ is $\sum_{i:(i,j)\in E}s_{m,ij}$; $N_{p}$ contains the top-rr% nodes by this value.
@@
Unless stated otherwise, in our implementation, we use k=5k{=}5 and a $r=10\%$ node fraction. This metric quantifies the degree of consensus among the most causally important heads i.e. measure top-k-head agreement on top-$r\%$ nodes. A high Focus score indicates that the model’s key components agree on the same salient graph substructures, simplifying interpretation.
@@ III-C5 Interpretation Guidelines
The combination of specialization and focus determines interpretability:
@@
  • High Spec + Low Focus: Important heads identifiable but attend to different nodes, suggesting complementary strategies.

### v2

@@ I Introduction
These discovered architectures must also be interpretable for trustworthiness and scientific validity. Unlike well-developed GNN explainability methods [21] that identify prediction-critical subgraphs, GT interpretation, widely regarded as a blackbox [22], remains dominated by attention visualization [23] lacking quantitative metrics and the ability to differentiate visually prominent but functionally irrelevant patterns from subtle yet critical ones. More recently, noteworthy efforts in GT interpretability have focused on analyzing attention graphs (i.e. aggregating head/layer attention into an attention graph) [24]. However, none of the prior methods test faithfulness (whether heads/nodes causally affect predictions) nor enable cross-model comparison via standardized scores. Task-specific self-explanations exist but do not generalize to arbitrary GTs or provide head attributions [25]. We seek a GT-agnostic instance-level interpretability framework grounded in perturbation/ablation that delivers model-level scores for comparing GT interpretability.
@@
Second, we introduce a Fidelity like [29, 30], instance-level, ablation-grounded score: Head-deviation, quantifying prediction changes when specific heads are masked. From per-head effects on each instance, we compute two complementary metrics: Specialization captures head distinguishability as dispersion of head impacts (high = few heads carry causal load; low = heads act interchangeably); Focus captures consensus over graph regions by measuring top-head agreement on top-k nodes (high = heads converge on same salient substructures or nodes). Aggregating per-instance values by dataset median yields robust model-level summaries. Together, these form our interpretability framework that we propose to analyse our model along with existing models for comparison. Head-deviation provides head and graph-region attributions per input, while median Specialization and median Focus enable dataset-level GT comparison indicating which model is easier to interpret at matched accuracy.
@@ III-A3 Feed-Forward Layers and Skip Connections
where $\sigma$ is the nonlinearity (we use ReLU in our implementation), $\mathbf{W}_{1}^{(\ell)}\in\mathbb{R}^{d\times rd}$, $\mathbf{W}_{2}^{(\ell)}\in\mathbb{R}^{rd\times d}$, and $\mathbf{b}_{1}^{(\ell)},\mathbf{b}_{2}^{(\ell)}$ are biases. We use r=2r{=}2 in our implementation, and $\mathrm{Norm}$ denotes either LayerNorm or BatchNorm depending on configuration.
@@ III-B DARTS-GT
where $\mathcal{S}$ denotes the search space of all $n^{L}$ architectures, $w$ represents the network weights, and $w^{*}(\mathcal{A})$ are the optimal weights for architecture $\mathcal{A}$. The following sections describe the approach we chose to solve the architecture search problem and the set of candidate operations, and hence, the search space.
@@ III-B1 Bilevel Optimization via DARTS
Enumerating all possible configurations in $\mathcal{S}$ is generally infeasible, so we adopt the Differential Architecture Search (DARTS) [27] framework to solve the optimization problem formulated by equations 11 and 12. DARTS reformulates discrete selection as a continuous, differentiable optimization. Specifically, it introduces a softmax relaxation over candidate operations, casting the architecture search as the following bilevel optimisation problem [38]:
@@
where $w$ represents the network weights and $\alpha=\{\alpha_{o}^{(\ell)}\mid o\in\mathcal{O},\ell\in\{1,...,L\}\}$ are continuous architecture parameters, with $\alpha_{o}^{(\ell)}\in\mathbb{R}$ representing the architecture weight (importance factor) for the operation $o$ at layer $\ell$. Rather than solving the full bilevel optimization, we adopt the efficient first-order approximation of DARTS [27], which avoids Hessian–vector products. In practice, we first randomly shuffle the entire training dataset $\mathcal{D}_{\text{train}}$ and partition it into darts-training split, $\mathcal{D}_{\text{DARTStrain}}$ for updating weights $w$ and darts-validation split, $\mathcal{D}_{\text{DARTSval}}$ for updating architecture parameters $\alpha$, alternating the single gradient steps [27]. More details about this approach can be found in [27]. Although the second-order DARTS update is more accurate, it requires expensive Hessian–vector products and offers limited practical gains in our setting.
@@ III-B2 Search Space Definition
In our work, we use the following set of candidate operations: $\mathcal{O}$ = {GINE [39], GATv2 [40], GatedGCN [32]}. Hence, $|\mathcal{O}|=n=3$ leads to a search space of $3^{L}$ architectures.
The following operators were selected because they represent very distinct message-passing paradigms that capture complementary aspects of graph structure.
@@ III-B3 Asymmetric Attention Design
Departing from the standard transformer formalism outlined in the Preliminaries section, the novelty of our approach includes an asymmetric design: queries are derived directly from node features, while keys and values are produced by layer-specific GNN operators. This is initially motivated by the GMH design of GMT [26], where they demonstrated that decoupling the computational sources for the attention components enhances graph-aware representations. In GMT, node features are first transformed by a GNN to obtain $H$, with queries taken directly as $Q=H$ and each head computing its own keys and values via an additional GNN applied to $H$. This asymmetry separates structural encoding from feature representation, but at the cost of computational inefficiency, since every key and value inside every head requires an independent GNN. In contrast, our design takes node features directly into $Q$, and shares the GNN operator across heads within a layer, producing an efficient formulation that still preserves the benefits of feature–structure decoupling. More broadly, rather than augmenting Graph Transformers externally with auxiliary MPNNs, a widely practiced approach [6, 16], we demonstrate that graph awareness can emerge from an architectural redesign of the attention mechanism itself.
@@
During the search phase, DARTS [27] relaxes the selection of discrete operators into a continuous mixture. At each layer $\ell$, instead of selecting a single operator, we compute a weighted combination:
@@
During search, the mixed operator output $\overline{o}^{(\ell)}(\mathbf{Z}^{(\ell)},E)$ provides the structural path for the keys and values:
@@
We compute $\overline{o}^{(\ell)}(\mathbf{Z}^{(\ell)},E)$ once per layer and share it across all heads to form $\mathbf{K}_{m}^{(\ell)}$ and $\mathbf{V}_{m}^{(\ell)}$, $m\in[1,\dots,M]$. After discretization, $\overline{o}^{(\ell)}$ is replaced by the selected operator $o^{(\ell)*}$(see 5) below). Attention may be dense or sparse, as that only changes the softmax normalization (all-pairs vs. edge-softmax) and does not alter the source of $\mathbf{K}_{m},\mathbf{V}_{m}$.
@@ III-B5 Discretization and Final Architecture Training
Following the standard DARTS protocol, the final discretized architecture $\mathcal{A}^{*}=\{o^{(1)*},...,o^{(L)*}\}$ is then re-trained from scratch using the full training data. In our implementation, we instantiate a fresh model containing only the selected operations per layer. The algorithm is discussed in Algorithm 1.
@@ III-B6 Edge Residual Stream
$\mathrm{MLP}_{\text{edge}}:\mathbb{R}^{d_{e}}\!\to\!\mathbb{R}^{d}$ is a two-layer MLP with dropout; $\sigma_{g}(x)$ is the sigmoid function, and $\gamma^{(\ell)}\!\in\!\mathbb{R}$ is a per-layer learnable scalar controlling the gate; $\mathbf{Y}^{(\ell)}$ is the pre-residual attention output. $\mathrm{MLP}_{\text{edge}}$ is shared across layers and $\mathbf{H}^{\text{edge}}$ is computed once per forward pass. Please see Supplementary Table I for ablation studies related to the edge residual stream. This edge residual stream operates independently of the selected GNN operations, allowing the model to leverage edge information even when the discretized architecture uses operations with limited utilization of edge features.
@@ III-B7 Computational Cost
Our proposed asymmetric block applies a single GNN per layer to produce the structural representations used in both $K$ and $V$. In contrast, GMT’s Graph Multi-head Attention (GMH) [26] instantiates per-head GNNs for both keys and values, i.e. $\{\mathrm{GNN}^{K}_{i},\mathrm{GNN}^{V}_{i}\}_{i=1}^{M}$. Consequently, the GNN forward cost in GMH scales as $\mathcal{O}(2M\cdot\text{cost}_{\text{GNN}})$, while in our design it is $\mathcal{O}(1\cdot\text{cost}_{\text{GNN}})$ per layer, where $\text{cost}_{\text{GNN}}$ denotes the computational cost of a single GNN forward pass per layer. Thus, we preserve the feature–structure interaction ($Q$ from features; $K,V$ from GNN) while reducing the structural computation across heads. From the perspective of architecture search overhead, DARTS [27] uses a first-order approximation that avoids Hessian–vector product overhead.
@@ III-C1 Framework Overview
We propose a systematic framework to quantify Graph Transformer interpretability through head-level and node-level analysis. Our approach goes beyond visualization and yields quantitative, instance-specific evidence of which components causally drive the predictions. We introduce Head-Deviation, Specialization, and Focus, aligned with established interpretability principles where fidelity, sparsity, and consensus are standard evaluation criteria [30, 29]. Because these metrics are ablation-based, they measure causal contribution rather than correlation, allowing robust, fair comparisons of interpretability across architectures.
@@
For head $m$ in layer $\ell$, we implement a weight-level ablation that removes its contribution at inference. Concretely, we silence head $m$ in layer $\ell$, by zeroing its pre-concatenation output $Y_{m}^{(\ell)}$ (equivalently, zeroing the corresponding slice of $\mathbf{W}_{\text{out}}$ in Eq. 6). All other parameters, including the outputs of other heads, layer norms, and feed-forward sublayers, are left unchanged. The models are evaluated in eval mode (no dropout), with a fixed seed and without any retraining or adaptation.
@@ III-C2 Head Deviation
To quantify the importance of individual attention heads, we measure how their removal affects the task loss. For a graph $G_{i}$ with the true label $y_{i}$, let $f(G_{i})$ denote the prediction of the model and $f^{-m,\ell}(G_{i})$ denote the prediction with head $m$ in layer $\ell$ masked.
@@
Here, $\mathrm{loss}(\cdot,\cdot)$ denotes the loss appropriate to the machine learning task (e.g., MAE for regression, binary cross-entropy for binary classification, cross-entropy for multi-class classification). We quantify head importance through causal intervention: the deviation $\delta_{m,\ell,i}$ measures how the prediction changes when head $m$ in layer $\ell$ is ablated, providing direct evidence of its functional necessity rather than mere correlational patterns. This approach asks ”what happens without this head?” , distinguishes truly important components from those that merely appear salient in visualizations.
@@
For each instance, we rank the heads by their deviation $\delta_{m,\ell,i}$ in descending order. Beneficial heads ($\delta_{m,\ell,i}>0$, masking increases loss) thus appear first, while harmful heads ($\delta_{m,\ell,i}<0$) appear last. When we need a sign-agnostic view, we report the magnitude $|\delta_{m,\ell,i}|$ as a measure of head impact.
@@ III-C3 Global Interpretability Metrics
where $\mathrm{std}(.)$ denotes the standard deviation across all heads’ deviations with $\mathcal{M}=\{1,...,M\}$, $\mathcal{L}=\{1,...,L\}$ where $M$ is the number of attention-heads and $L$ total number of layers, defined earlier.
High specialization indicates that heads are easily distinguishable by their causal impact, whereas low specialization suggests similar head impacts, making it harder to isolate drivers unless complemented by the focus metric.
@@
Let $\mathcal{P}_{i}^{*}=\{p_{1},\dots,p_{k}\}$ be the top-$k$ head-layer pairs for instance $G_{i}$, ranked according to the head-ranking rule defined above, where $p_{j}=(i^{\prime}_{j},i^{\prime\prime}_{j})$, j=1,,kj=1,...,k, $i^{\prime}_{j}\in\mathcal{M}$, $i^{\prime\prime}_{j}\in\mathcal{L}$ and $i^{\prime}_{j}$, $i^{\prime\prime}_{j}$ are index variables. For a pair $p$, we define $N_{p}$ as the set of top-rr% nodes by incoming attention mass for that head on $G_{i}$. $N_{p}$ contains the top-rr% nodes by the following values:
  • Dense attention: with head softmax matrix $\mathbf{S}^{(\ell)}_{m}$, incoming mass for node $j$ is $\sum_{i}[\mathbf{S}^{(\ell)}_{m}]_{ij}$;
  • Sparse attention: with edge-softmax $s_{m,ij}$ on $(i,j)\in E$, incoming mass for node $j$ is $\sum_{i:(i,j)\in E}s_{m,ij}$.
@@
Unless stated otherwise, in our implementation, we use k=5k{=}5 and a $r=10\%$ node fraction. This metric quantifies the degree of consensus among the most causally important heads, i.e. measure top-k-head agreement on top-$r\%$ nodes. A high Focus score indicates that the model’s key components agree on the same salient graph substructures, simplifying interpretation.
@@ III-C5 Interpretation Guidelines
The combination of specialization and focus determines how interpretable an instance-level prediction is.
@@
  • High Spec + Low Focus: Important heads are identifiable but attend to different nodes, suggesting complementary strategies.